OpenStudy (anonymous):

write 0.438 as a fraction in simplest form

OpenStudy (texaschic101):

do you know the place value of the 8 ?

OpenStudy (anonymous):

0.438=438/1000=219/500

OpenStudy (texaschic101):

nvm....you already got the answer given to you